# Splittings of One-Ended Groups with One-Ended Halfspaces

### v3

@@ 5 Groups $G$ that are simply connected at infinity; have $H^{2}(G,\mathbb{Z}G)=\{0\}$
Definition 13. The space $X$ is simply connected at infinity if for any compact set $C$ there is a compact set $D$ such that loops $\alpha:[0,1]\to X-D$ are homotopically trivial in $X-C$. This means there is a homotopy $H:[0,1]\times[0,1]\to X-C$ such that $H(t,0)=\alpha(t)$, $H(0,t)=H(1,t)=H(t,1)=H(0,0)$ for all $t\in[0,1]$.
Definition 14. The space $X$ is 1-acyclic at infinity (has pro-finite first homology at infinity) if for any compact set $C$ there is a compact set $D$ such that the image of $H_{1}(X-D)$ in $H_{1}(X-C)$ under the homomorphism induced by the inclusion of $X-D$ into $X-C$ is trivial (respectively, finite). A loop $\alpha$ in $X-C$ is homologically trivial if there is an orientable 2-manifold $M$ bounded by an embedded loop $\alpha^{\prime}$ and a continuous map $H:M\to X-C$ such that $H$ restricted to $\alpha^{\prime}$ maps (in the obvious way) to $\alpha$ (see Proposition 12.8, [GH81]).
@@
Assume that $X$ has pro-finite first homology at infinity. Let $X_{0}=X_{1}\cap X_{2}$. Let $D$ be a finite subcomplex of $X$ containing $K$ such that for any loop $\tau$ in $X-D$, there is $n>0$ such that $\tau^{n}$ is homologically trivial in $X-K$. Since both $X_{1}$ and $X_{2}$ are one-ended, there is a finite subcomplex $E$ of $X$ containing $D$ such that any two vertices of $X_{0}-E$ can be joined by an edge path in $X_{1}-D$ and by an edge path in $X_{2}-D$. Choose vertices $v$ and $w$ in $X_{0}-E$ such that $v$ and $w$ cannot be joined by an edge path in $X_{0}-K$ (so $v\in V$ and $w\in W$ where $V$ and $W$ are distinct components of $X_{0}-K$). Let $\alpha$ be an edge path in $X_{1}-D$ from $v$ to $w$ and let $\beta$ be an edge path in $X_{2}-D$ from $w$ to $v$. The concatenation, $(\alpha,\beta)$, is a loop in $X-D$. By our earlier assumption, $(\alpha,\beta)^{n}$ is homologically trivial in $X-K$. Let $M$ be an orientable 2-manifold bounded by an embedded loop $(\alpha^{\prime}_{1},\beta_{1}^{\prime},\ldots,\alpha_{n}^{\prime},\beta_{n}^{\prime})$ and let $H:M\to X-K$ be a continuous map such that $H$ restricted to $\alpha_{i}^{\prime}$ and $\beta_{i}^{\prime}$ maps (in the obvious way) to the edge paths $\alpha$ and $\beta$ respectively. Let $v_{i}^{\prime}$ and $w_{i}^{\prime}$ be the initial and terminal points of $\alpha_{i}^{\prime}$ (See Figure 3).
@@
The sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ are disjoint closed subsets of $M$ containing $\{w_{1}^{\prime},\ldots w_{m}^{\prime}\}^{\prime}$ and $\{v_{1}^{\prime},\ldots,v_{n}^{\prime}\}$ respectively.
@@
Consider the circle $S^{1}$ with diametrically opposite vertices $\bar{v}$ and $\bar{w}$. Let $T$ and $B$ be the two closed arcs in $S^{1}$ such that $T\cup B=S^{1}$ and $T\cap B=\{\bar{v},\bar{w}\}$. Map the (disjoint) closed sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ in $H^{-1}(X_{1})$ to $\{\bar{w}\}$ and $\{\bar{v}\}$ respectively. Extend this map by Tietze’s extension theorem to a continuous map $f_{1}:H^{-1}(X_{1})\to T$. Similarly define $f_{2}:H^{-1}(X_{2})\to B$ (with $f_{2}(H^{-1}(W))=\{\bar{w}\}$ and $f_{2}(H^{-1}(X_{0}-W))=\{\bar{v}\}$). Since $f_{1}$ and $f_{2}$ agree on the closed set $H^{-1}(X_{1})\cap H^{-1}(X_{2})=H^{-1}(X_{0})=H^{-1}(W)\cup H^{-1}(X_{0}-W)$, we have a continuous function $f:M\to S^{1}$ that agrees with $f_{1}$ and $f_{2}$ on $H^{-1}(X_{1})$ and $H^{-1}(X_{2})$ respectively. Note that for each $i$, $f\alpha_{i}^{\prime}=f_{1}\alpha_{i}^{\prime}=\bar{\alpha}_{i}$ is a path in $T$ from $\bar{v}$ to $\bar{w}$ and $f\beta_{i}^{\prime}=f_{2}\beta_{i}^{\prime}=\bar{\beta}_{i}$ is a path in $B$ from $\bar{w}$ to $\bar{v}$. Furthermore, for each $i$, the homotopy class of loop $(\bar{\alpha}_{i},\bar{\beta}_{i})$ at $\bar{v}$ generates $\pi_{1}(S_{1},\bar{v})=\mathbb{Z}$. The contradiction arises since the homotopy class of the loop $(\alpha_{1}^{\prime},\beta_{1}^{\prime},\ldots,\alpha_{n}^{\prime},\beta_{n}^{\prime})$ is a commutator in $\pi_{1}(M,v^{\prime})$ and so maps to the trivial element under the homomorphism $f_{\ast}:\pi_{1}(M,v^{\prime})\to\pi_{1}(S^{1},\bar{v})$. But instead this class is mapped by $f_{\ast}$ to the class of $(\bar{\alpha}_{1},\bar{\beta}_{1},\ldots,\bar{\alpha}_{n},\bar{\beta}_{n})$, a non-trivial element of $\pi_{1}(S^{1},\bar{v})$. ∎
@@
7 Artificial splittings and an example
We give an example of a finitely presented one-ended simply connected at infinity group that splits non-trivially over a finitely generated infinite-ended group. So there is no result that implies a general non-trivial splitting of a one-ended finitely presented group over an infinite-ended group is not simply connected at infinity. The next lemma describes how a splitting of a group can “artificially” induce a splitting over a larger subgroup.
@@ 7 Artificial splittings and an example
Let $\mathcal{P}_{Q}=\langle S_{Q}:R_{Q}\rangle$ be a presentation for $Q$, where $Q\in\{A,B,C,D\}$. Assume that $\mathcal{P}_{C}$ is a sub-presentation of $\mathcal{P}_{A}$ and $\mathcal{P}_{D}$ (so that $S_{C}\subset S_{A}$, $R_{C}\subset R_{A}$, $S_{C}\subset S_{D}$ and $R_{C}\subset R_{D}$). Also assume that $\mathcal{P}_{D}$ is a sub-presentation of $\mathcal{P}_{B}$. Then a presentation for $G$ is
@@
Let $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. The group $G$ artificially decomposes as
$G=(A\ast_{C}D)\ast_{D}B$
so the one-ended simply connected at infinity group $G$ splits non-trivially over the infinite-ended group $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. Note that $A\ast_{C}D=A\ast\mathbb{Z}_{c}$ so we also have
$G=(A\ast\mathbb{Z}_{c})\ast_{D}B.$
Hence, in order to recognize this last splitting as an artificial splitting of $G$ we would have to recognize that $A\ast\mathbb{Z}_{c}$ splits non-trivially with $D$ as a vertex group. Observe that $A\ast_{C}D$ is an artificial splitting of $A\ast\mathbb{Z}_{c}$. Note also that the halfspace corresponding to $A*\mathbb{Z}_{c}$ in the last splitting has more than one end by Corollary 1.5.
@@ 8 Some examples of splittings with one-ended halfspaces
(Co)homological dimension: Any nontrivial amalgam $G=A*_{C}B$ or HNN extension $G=A*_{C}$ of an $n$-dimensional duality group $G$ requires $n-1\leq\mathrm{cd}(C)\leq\mathrm{cd}(A),\mathrm{cd}(B)\leq n$ [Bie81, Proposition 9.14]. Furthermore, if $G$ is a Poincaré duality group, then every subgroup of $G$ either has finite index or has homology dimension at most $n-1$ [Bie81, Proposition 9.22].
@@
Note that the base group $A\ast\langle s\rangle$ has cohomological dimension $n$ and the edge group $C$ has cohomological dimension $n-1$, so $G$ looks like a duality group from the perspective of cohomological dimension. The point of including the element $as_{1}sa_{2}\cdots sa_{n}s$ in $C$ is to prevent other splittings of $G$ over groups of lower cohomological dimension.

### v4

@@ 5 Groups $G$ that are simply connected at infinity; have $H^{2}(G,\mathbb{Z}G)=\{0\}$
Definition 13. The space $X$ is simply connected at infinity if for any compact set $C$ there is a compact set $D$ such that loops $\alpha:[0,1]\to X-D$ are homotopically trivial in XCX-C. This means there is a homotopy $H:[0,1]\times[0,1]\to X-C$ such that $H(t,0)=\alpha(t)$, $H(0,t)=H(1,t)=H(t,1)=H(0,0)$ for all $t\in[0,1]$.
Definition 14. The space $X$ is 1-acyclic at infinity (has pro-finite first homology at infinity) if for any compact set $C$ there is a compact set $D$ such that the image of $H_{1}(X-D)$ in $H_{1}(X-C)$ under the homomorphism induced by the inclusion of XDX-D into XCX-C is trivial (respectively, finite). A loop $\alpha$ in XCX-C is homologically trivial if there is an orientable 2-manifold $M$ bounded by an embedded loop $\alpha^{\prime}$ and a continuous map $H:M\to X-C$ such that $H$ restricted to $\alpha^{\prime}$ maps (in the obvious way) to $\alpha$ (see Proposition 12.8, [GH81]).
@@
Assume for contradiction that $X$ has pro-finite first homology at infinity. Let $X_{0}=X_{1}\cap X_{2}$. Let $D$ be a finite subcomplex of $X$ containing $K$ such that for any loop $\tau$ in XDX-D, there is $n>0$ such that $\tau^{n}$ is homologically trivial in XKX-K. Since both $X_{1}$ and $X_{2}$ are one-ended, there is a finite subcomplex $E$ of $X$ containing $D$ such that any two vertices of $X_{0}-E$ can be joined by an edge path in $X_{1}-D$ and by an edge path in $X_{2}-D$. Choose vertices $v$ and $w$ in $X_{0}-E$ such that $v$ and $w$ cannot be joined by an edge path in $X_{0}-K$ (so $v\in V$ and $w\in W$ where $V$ and $W$ are distinct components of $X_{0}-K$). Let $\alpha$ be an edge path in $X_{1}-D$ from $v$ to $w$ and let $\beta$ be an edge path in $X_{2}-D$ from $w$ to $v$. The concatenation, $\alpha*\beta$, is a loop in XDX-D. By our earlier assumption, $(\alpha*\beta)^{n}$ is homologically trivial in XKX-K. Let $M$ be an orientable 2-manifold bounded by an embedded loop $\alpha^{\prime}_{1}*\beta_{1}^{\prime}*\ldots*\alpha_{n}^{\prime}*\beta_{n}^{\prime}$ and let $H:M\to X-K$ be a continuous map such that $H$ restricted to $\alpha_{i}^{\prime}$ and $\beta_{i}^{\prime}$ maps (in the obvious way) to the edge paths $\alpha$ and $\beta$ respectively. Let $v_{i}^{\prime}$ and $w_{i}^{\prime}$ be the initial and terminal points of $\alpha_{i}^{\prime}$ (See Figure 4).
@@
The sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ are disjoint closed subsets of $M$ containing $\{w_{1}^{\prime},\ldots,w_{m}^{\prime}\}$ and $\{v_{1}^{\prime},\ldots,v_{n}^{\prime}\}$ respectively.
@@
Consider the circle $S^{1}$ with diametrically opposite vertices $\bar{v}$ and $\bar{w}$. Let $T$ and $B$ be the two closed arcs in $S^{1}$ such that $T\cup B=S^{1}$ and $T\cap B=\{\bar{v},\bar{w}\}$. Map the (disjoint) closed sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ in $H^{-1}(X_{1})$ to $\{\bar{w}\}$ and $\{\bar{v}\}$ respectively. Extend this map by Tietze’s extension theorem to a continuous map $f_{1}:H^{-1}(X_{1})\to T$. Similarly define $f_{2}:H^{-1}(X_{2})\to B$ (with $f_{2}(H^{-1}(W))=\{\bar{w}\}$ and $f_{2}(H^{-1}(X_{0}-W))=\{\bar{v}\}$). Since $f_{1}$ and $f_{2}$ agree on the closed set $H^{-1}(X_{1})\cap H^{-1}(X_{2})=H^{-1}(X_{0})=H^{-1}(W)\cup H^{-1}(X_{0}-W)$, we have a continuous function $f:M\to S^{1}$ that agrees with $f_{1}$ and $f_{2}$ on $H^{-1}(X_{1})$ and $H^{-1}(X_{2})$ respectively. Note that for each $i$, $f\alpha_{i}^{\prime}=f_{1}\alpha_{i}^{\prime}=\bar{\alpha}_{i}$ is a path in $T$ from $\bar{v}$ to $\bar{w}$ and $f\beta_{i}^{\prime}=f_{2}\beta_{i}^{\prime}=\bar{\beta}_{i}$ is a path in $B$ from $\bar{w}$ to $\bar{v}$. Furthermore, for each $i$, the homotopy class of loop $(\bar{\alpha}_{i},\bar{\beta}_{i})$ at $\bar{v}$ generates $\pi_{1}(S_{1},\bar{v})=\mathbb{Z}$. The contradiction arises since the homotopy class of the loop $\alpha_{1}^{\prime}*\beta_{1}^{\prime}*\ldots*\alpha_{n}^{\prime}*\beta_{n}^{\prime}$ is a commutator in $\pi_{1}(M,v^{\prime})$ and so maps to the trivial element under the homomorphism $f_{\ast}:\pi_{1}(M,v^{\prime})\to\pi_{1}(S^{1},\bar{v})$. But instead this class is mapped by $f_{\ast}$ to the class of $\bar{\alpha}_{1}*\bar{\beta}_{1}*\ldots*\bar{\alpha}_{n}*\bar{\beta}_{n}$, a non-trivial element of $\pi_{1}(S^{1},\bar{v})$. ∎
@@
7 An example of a splitting obtained from folding
We give an example of a finitely presented one-ended simply connected at infinity group that splits non-trivially over a finitely generated infinite-ended group. So there is no result that implies a general non-trivial splitting of a one-ended finitely presented group over an infinite-ended group is not simply connected at infinity. Our example is obtained via folding, an operation originally introduced for graphs by Stallings [Sta83], and applied to group splittings by Bestvina and Feighn [BF91]. More precisely, our example is obtained via the following lemma.
@@ 7 An example of a splitting obtained from folding
We can transform the Bass–Serre tree of the first splitting into the Bass–Serre tree of the second splitting by folding together the edges that correspond to cosets of $C$ which lie in the same coset of $D$. If $D$ is finitely generated, then this can be expressed as a finite sequence of folds in the sense of Bestvina–Feighn [BF91].
One can also argue using group presentations as follows. Let $\mathcal{P}_{Q}=\langle S_{Q}:R_{Q}\rangle$ be a presentation for $Q$, where $Q\in\{A,B,C,D\}$. Assume that $\mathcal{P}_{C}$ is a sub-presentation of $\mathcal{P}_{A}$ and $\mathcal{P}_{D}$ (so that $S_{C}\subset S_{A}$, $R_{C}\subset R_{A}$, $S_{C}\subset S_{D}$ and $R_{C}\subset R_{D}$). Also assume that $\mathcal{P}_{D}$ is a sub-presentation of $\mathcal{P}_{B}$. Then a presentation for $G$ is
@@
Let $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. As in Lemma 7.1, the group $G$ has a second decomposition as
$G=(A\ast_{C}D)\ast_{D}B,$
so the one-ended simply connected at infinity group $G$ splits non-trivially over the infinite-ended group $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. Note also that $B$ is one-ended, so the halfspace corresponding to $B$ in the above splitting is one-ended by Lemma 2.9, however, $D$ is infinite-ended, so the halfspace corresponding to $A\ast_{C}D$ has more than one end by Corollary 1.5.
@@ 8 Some examples of splittings with one-ended halfspaces
(Co)homological dimension: Any nontrivial amalgam $G=A*_{C}B$ or HNN extension $G=A*_{C}$ of an $n$-dimensional duality group $G$ requires $n-1\leq\mathrm{cd}(C)\leq\mathrm{cd}(A),\mathrm{cd}(B)\leq n$ [Bie81, Proposition 9.14]. Furthermore, if $G$ is a Poincaré duality group, then every subgroup of $G$ either has finite index or has homology dimension at most n1n-1 [Bie81, Proposition 9.22].
@@
Note that the base group $A\ast\langle s\rangle$ has cohomological dimension $n$ and the edge group $C$ has cohomological dimension n1n-1, so $G$ looks like a duality group from the perspective of cohomological dimension. The point of including the element $as_{1}sa_{2}\cdots sa_{n}s$ in $C$ is to prevent other splittings of $G$ over groups of lower cohomological dimension.